# thanks: M.C. and D.P. contributed equally to this workthanks: M.C. and D.P. contributed equally to this work

### v2

@@ IV Results
Figure 10: As in Fig. 7, but for primordial regular Hayward-like black holes, with values of the regularizing parameter $\ell/r_{H}=0.15$ (red dotted curve), $0.3$ (green dashed curve), $0.45$ (magenta dash-dotted curve), $0.6$ (black dash-dotdotted curve), and $0.75$ (brown dash-dotdotdotted curve).
@@
For each of the PRBHs discussed in Sec. II, we now proceed to derive upper limits on $f_{\text{pbh}}$ as a function of the PRBH mass $M_{\text{pbh}}$, for different values of $\ell$, using the methodology presented in Sec. III.3. The results are shown in Figs. 7,  8,  9, and 10 for the Bardeen, Hayward, CGSV, and Hayward-like PRBHs respectively. For each case, we also plot the constraints on $f_{\text{pbh}}$ for the $\ell=0$ case (blue solid curve in all the Figures), which correspond to the standard Schwarzschild PBH scenario widely studied in the literature. As a sanity check, we have verified that our $\ell=0$ constraints exactly recover those of the seminal Ref. Carr et al. (2010). It is worth noting that, for any given value of $\ell$, the value of $M_{\text{pbh}}$ corresponding to the upper right edge of the $f_{\text{pbh}}$ constraints (i.e. the value of $M_{\text{pbh}}$ for which the limit reads $f_{\text{pbh}}<1$) marks the lower edge of the (modified – either enlarged or contracted) asteroid mass window.
@@
We begin by discussing the Bardeen, Hayward, and CGSV PRBHs, for which we saw earlier that the temperature and photon spectra decrease in intensity with increasing regularizing parameter $\ell$ (see the discussion in Sec. III.2, and Figs. 25). As we could have expected, this behaviour leads to overall looser constraints on $f_{\text{pbh}}$ (for any given $M_{\text{pbh}}$) relative to the standard limits reported for Schwarzschild PBHs in the literature. In the case of near-extremal Hayward PRBHs ($\ell=0.45r_{H}$) this behaviour is somewhat enhanced compared to the near-extremal Bardeen and CGSV PRBHs, with the upper limits on $f_{\text{pbh}}$ approximately three orders of magnitude looser than the corresponding Schwarzschild ones: again, this is somewhat unsurprising when comparing Fig. 3 to Figs. 2 and 4. This could also have been expected from Fig. 1, noting that the temperature of Hayward BHs decreases more rapidly with increasing $\ell/r_{H}$ relative to the Bardeen and especially CGSV ones. Although the temperature is not the only factor at play in determining the resulting evaporation spectra, given that the GBFs also play a key role as per Eq. (29), it is reassuring to see that the temperature behaviour observed in Fig. 1 is qualitatively reflected in the limits on $f_{\text{pbh}}$ we derive.
For the Hayward-like BH, we observe exactly the opposite trend, with overall tighter constraints on $f_{\text{pbh}}$ relative to the Schwarzschild ones. Again, such a behaviour is in line with expectations given the behaviour of the temperature observed in Fig. 1, and of the spectra as shown in Fig. 5. The relative shift in $f_{\text{pbh}}$ constraints for a given value of $M_{\text{pbh}}$ and $\ell/r_{H}$ is somewhat less dramatic than what we observed for the Bardeen, Hayward, and CGSV BHs, with shifts of less than two orders of magnitude.
@@
As a result of the shifts discussed above, the lower edge of the asteroid mass window where PBHs could make up the entire DM component is modified for all four metrics considered. We recall that in the Schwarzschild case, the lower edge of the window lies at $M_{\text{pbh}}\simeq 10^{14}\,{\text{kg}}$. For the Bardeen, Hayward, and CGSV PRBHs, the looser constraints on $f_{\text{pbh}}$ result in the asteroid mass window further opening up by up to half a decade in mass or more. The maximum extension of the window is reached for the Hayward PRBH closer to extremality, in which case the lower edge decreases by about an order of magnitude to $M_{\text{pbh}}\simeq 10^{13}\,{\text{kg}}$. The opposite behaviour is of course observed for Hayward-like PRBHs, in which case the asteroid mass window further closes down, although by a lesser extent with respect to the previous cases. For instance, for $\ell/r_{H}=0.75$, the lower edge increases to $M_{\text{pbh}}\simeq 2\times 10^{14}\,{\text{kg}}$. Overall, we therefore observe that considering PRBHs in place of the standard Schwarzschild ones can move the resulting constraints on $f_{\text{pbh}}$ in either direction, further opening up or closing down the asteroid mass window, with the allowed region for the window lower edge spanning over a decade in mass, at least for the PRBHs considered.
@@
Three comments are in order before concluding. Firstly we note that, for a given PRBH space-time, the curves describing the $f_{\text{pbh}}(M)$ limits are approximately, but not exactly parallel to the Schwarzschild ones (blue solid curves in Figs. 710). The reason is simply that, as $\ell$ is increased, the datapoint shown in Fig. 6 which is first being overshot and therefore responsible for determining the $f_{\text{pbh}}$ limit can potentially change (in part due to the spectrum slightly changing shape, especially in the Hayward-like case where we saw that an increase in $\ell$ lead to a richer peak structure due to the effect of different $l$ modes).
@@
Next, the constraints we have determined on $f_{\text{pbh}}$ at a fixed value of $\ell/r_{H}$ implicitly assume that all PRBHs in the Universe carry the same value of “hair” parameter $\ell$. However, particularly given our agnostic stance with regards to the origin of these space-times, in principle the value of $\ell/r_{H}$ can vary from PRBH to PRBH. To make an analogy, let us assume for a moment that Reissner-Nordström BHs are astrophysically relevant. Then, since the electric charge $Q$ is not tied to a universal parameter of the underlying Einstein-Maxwell Lagrangian, there is no reason to expect it to carry the same value across all BHs. In the language of Ref. Vagnozzi et al. (2023), the regularizing parameter for all four PRBHs considered is a “specific hair” rather than an “universal hair” (see Ref. Vagnozzi et al. (2023) for various examples of BH solutions carrying universal hair), unless one were able to tie $\ell$ to some fundamental parameter of the underlying theory, which however is not the case in the phenomenological approach we are following. In principle one should therefore account for the (non-monochromatic) $\ell$ distribution for PRBHs across the Universe to determine constraints on $f_{\text{pbh}}$. We see no obvious way of doing this, while noting that such a procedure would most likely result in upper limits on $f_{\text{pbh}}$ lying between the Schwarzschild and extremal cases: this observation suffices for our pilot study, and we defer a more complete investigation to future work.
@@
Our final comment concerns the fact that evaporation limits on the PBH abundance are not the only ones at play. Indeed, as recently summarized in Ref. Carr and Kuhnel (2022), there are essentially four classes of limits, each of which is relevant in a different mass range: evaporation, lensing, dynamical, and accretion constraints. Constraints from the accretion of background gas at early times are relevant in a completely different mass range ($10^{30}\lesssim M_{\text{pbh}}/{\text{kg}}\lesssim 10^{37}$ – see Fig. 7 of Ref. Carr and Kuhnel (2022) and Fig. 10 of Ref. Carr et al. (2021d)). Although these have been derived assuming Schwarzschild PBHs, moving to the PRBH picture we have considered will not shift the relevant mass range by the $\gtrsim 18$ orders of magnitude required for these constraints to compete with the evaporation ones, unless the physics of gas accreting around RBHs changes drastically with respect to the standard picture, which appears very unlikely. Dynamical constraints, most of which are associated to the destruction of different astronomical objects by the passage of nearby PBHs, are also relevant in a completely different mass range ($10^{31}\lesssim M_{\text{pbh}}/{\text{kg}}\lesssim 10^{52}$ – see Fig. 7 of Ref. Carr and Kuhnel (2022) and Fig. 10 of Ref. Carr et al. (2021d)), and considerations completely analogous to those we made for accretion constraints hold.  [^7] Potential exceptions to this mass range for dynamical constraints are those from capture of PBHs by white dwarfs or neutron stars at the centres of globular clusters Capela et al. (2013a, b); Pani and Loeb (2014), or from supernovae explosions resulting from transit of a PBH through a white dwarf Graham et al. (2015). However, these limits are highly disputed because of uncertainties in the dark matter density in globular clusters Ibata et al. (2013); Defillon et al. (2014), or based on the results of hydrodynamical simulations Montero-Camacho et al. (2019). For these reason, we will not consider the previously mentioned limits in our discussion.
@@
Of potentially more relevance to the present work are lensing constraints, which constrain the abundance of PBHs (and more generally massive compact halo objects) with masses $M_{\text{pbh}}\gtrsim 10^{12}\,{\text{kg}}$. Indeed, it is lensing constraints which locate the upper edge of the asteroid mass window where PBHs can make up all the DM. Nevertheless, we expect that these constraints should not change when moving from the Schwarzschild PBH framework to the PRBHs considered in this work. Indeed, with all other quantities being fixed (mass of source, relative distances, and so on), lensing constraints only depend on the lens mass $M$, and are unaffected by the metric structure of the lens. Therefore, at fixed mass $M$, we can assume that the lensing limits on Schwarzschild PBHs hold for our PRBHs as well. Note that, as already pointed out in footnote 2, the parameter $M$ appearing in the RBH metrics can be unambiguously identified with the RBH mass, just as with the parameter $M$ in the Schwarzschild metric. We can therefore conclude that for the PRBHs we are considering it should only be the lower edge of the asteroid mass window which is altered with respect to the Schwarzschild case, but not the upper edge. In other words, space-times for which the lower edge moves towards lower masses (as in the Bardeen, Hayward, and CGSV PRBH cases) genuinely correspond to an enlarged asteroid mass window, and conversely for the Hayward-like case where the lower edge moves towards higher masses. Therefore, the window where Bardeen, Hayward, and CGSV PRBHs could account for all the DM is much larger compared to that of Schwarzschild PBHs.
@@ V Conclusions
Over the past decade, primordial black holes have regained tremendous interest as viable dark matter candidates, with the so-called “asteroid mass window” ($10^{14}\,{\text{kg}}\lesssim M_{\text{pbh}}\lesssim 10^{20}\,{\text{kg}}$) where PBHs could potentially account for the entire DM currently still open. Nearly all works on PBHs assume that these are Schwarzschild or Kerr BHs. However, while phenomenologically perfectly valid, such an assumption may stir some unease on the theoretical side, due to the appearance of singularities in these metrics. In our work, we have conducted a pilot study aimed at addressing a question which naturally merges the DM and singularity problems, arguably two among the most important open problems in theoretical physics: “What if PBHs are non-singular”? Our study of primordial regular BHs (PRBHs) has focused on four so-called tr-symmetric metrics (including the well-known Bardeen and Hayward space-times), whereas our companion paper Calzà et al. (2024a) considers non-tr-symmetric metrics, including various metrics inspired from loop quantum gravity.
@@
We show that evaporation constraints on $f_{\text{pbh}}$, the fraction of DM in the form of PRBHs, can be substantially altered in either direction when moving away from the Schwarzschild picture, leading to the asteroid mass window further opening up or closing down depending on the direction of the shifts in $f_{\text{pbh}}$ limits. For three of the PRBHs we considered (the Bardeen, Hayward, and CGSV ones) the lower edge of the asteroid mass window is shifted by nearly a decade in mass, leading to a larger region of parameter space where PRBHs could account for the entire DM component, which should be the target of the same probes proposed to test the standard window Ray et al. (2021); Ghosh and Mishra (2024); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2023); Dent et al. (2024); Tamta et al. (2024). The opposite trend takes place with the Hayward-like BH we constructed, and we have argued that part of this different behaviour can be traced back to the evolution (increase or decrease) of the PRBH temperature as the regularizing parameter is increased. On the other hand, the nature of the regular BH core (de Sitter or Minkowski) does not appear to play a significant role in this sense. Overall, we have shown that the phenomenology of primordial regular BHs can be particularly rich, making the associated simultaneous solution to the DM and singularity problems one worthy of further study.
@@
We remark that the present work (alongside our companion paper Calzà et al. (2024a)) should be intended as a pilot study, and there are a huge number of interesting follow-up directions. One interesting avenue for further work involves systematically revisiting other sources of constraints which have been studied in the Schwarzschild PBH case, including but not limited to lensing, accretion, and dynamical constraints: while we have argued that these should not alter our considerations on the asteroid mass window, a detailed study which would allow us to extend our constraints over a much larger region of $M_{\text{pbh}}$-$f_{\text{pbh}}$ plane is nevertheless in order. In addition, the metrics we have considered are inherently phenomenological in nature, and it would therefore be worth extending our study to non-singular metrics which enjoy a strong theoretical motivation (our companion paper Calzà et al. (2024a) goes partially in this direction), including potentially metrics which are coupled to the cosmological expansion Farrah et al. (2023); Cadoni et al. (2023, 2024); Faraoni and Rinaldi (2024); Calzà et al. (2024b). Last but definitely not least, if PBHs are truly regular, one would hope to ascertain this via signatures complementary to those we have studied (for instance through gravitational wave signatures, VLBI imaging, particle motion, or energy extraction). We plan to address these and other related points in follow-up work.

### v3

@@ IV Results
For each of the PRBHs discussed in Sec. II, we now proceed to derive upper limits on $f_{\text{pbh}}$ as a function of the PRBH mass $M_{\text{pbh}}$, for different values of $\ell$, using the methodology presented in Sec. III.3. The results are shown in Figs. 7,  8, and 9 for Bardeen, Hayward, and CGSV PRBHs respectively. For each case, we also plot the constraints on $f_{\text{pbh}}$ for the $\ell=0$ case (blue solid curve in all the Figures), which correspond to the standard Schwarzschild PBH scenario widely studied in the literature. As a sanity check, we have verified that our $\ell=0$ constraints exactly recover those of the seminal Ref. Carr et al. (2010). It is worth noting that, for any given value of $\ell$, the value of $M_{\text{pbh}}$ corresponding to the upper right edge of the $f_{\text{pbh}}$ constraints (i.e. the value of $M_{\text{pbh}}$ for which the limit reads $f_{\text{pbh}}<1$) marks the lower edge of the asteroid mass window.
@@
For all PRBHs, we saw earlier that the temperature and photon spectra decrease in intensity with increasing regularizing parameter $\ell$ (see the discussion in Sec. III.2, and Figs. 35). As we could have expected, this behaviour leads to overall looser constraints on $f_{\text{pbh}}$ (for any given $M_{\text{pbh}}$) relative to the standard limits reported for Schwarzschild PBHs in the literature. In the case of near-extremal Hayward PRBHs ($\ell=0.45r_{H}$) this behaviour is somewhat enhanced compared to the near-extremal Bardeen and CGSV PRBHs, with the upper limits on $f_{\text{pbh}}$ approximately three orders of magnitude looser than the corresponding Schwarzschild ones: again, this is somewhat unsurprising when comparing Fig. 4 to Figs. 3 and 5. This could also have been expected from Fig. 1, noting that the temperature of Hayward BHs decreases more rapidly with increasing $\ell/r_{H}$ relative to the Bardeen and especially CGSV ones. Although the temperature is not the only factor at play in determining the resulting evaporation spectra, given that the GBFs also play a key role as per Eq. (27), it is reassuring to see that the temperature behaviour observed in Fig. 1 is qualitatively reflected in the limits on $f_{\text{pbh}}$ we derive.
@@
As a result of the shifts discussed above, the lower edge of the asteroid mass window where PBHs could make up the entire DM component is modified for all three metrics considered. We recall that in the Schwarzschild case, the lower edge of the window lies at $M_{\text{pbh}}\simeq 10^{17}\,{\text{g}}$. For the PRBHs we consider, the looser constraints on $f_{\text{pbh}}$ result in the asteroid mass window further opening up by up to half a decade in mass or more. The maximum extension of the window is reached for the Hayward PRBH closer to extremality, in which case the lower edge decreases by about an order of magnitude to $M_{\text{pbh}}\simeq 10^{16}\,{\text{g}}$. Overall, we therefore observe that considering PRBHs in place of the standard Schwarzschild ones can relax the resulting constraints on $f_{\text{pbh}}$, further opening up the asteroid mass window. The allowed region for the window lower edge spans over a decade in mass, at least for the PRBHs and range of $\ell$ considered. We note that, as $\ell$ moves towards the extremal limit, the temperatures of these PRBHs $T$ approaches zero. This is indeed generally expected from thermodynamical arguments. In this case, which is the one studied in Ref. Pacheco (2018), the PRBHs do not evaporate, and therefore our evaporation constraints do not apply. See also Refs. Davies et al. (2024) for a related study appearing after ours.
@@
Three comments are in order before concluding. Firstly we note that, for a given PRBH space-time, the curves describing the $f_{\text{pbh}}(M)$ limits are approximately, but not exactly parallel to the Schwarzschild ones (blue solid curves in Figs. 79). The reason is simply that, as $\ell$ is increased, the datapoint shown in Fig. 6 which is first being overshot and therefore responsible for determining the $f_{\text{pbh}}$ limit can potentially change (in part due to the spectrum slightly changing shape).
@@
Next, the constraints we have determined on $f_{\text{pbh}}$ at a fixed value of $\ell/r_{H}$ implicitly assume that all PRBHs in the Universe carry the same value of “hair” parameter $\ell$. However, particularly given our agnostic stance with regards to the origin of these space-times, in principle the value of $\ell/r_{H}$ can vary from PRBH to PRBH. To make an analogy, let us assume for a moment that Reissner-Nordström BHs are astrophysically relevant. Then, since the electric charge $Q$ is not tied to a universal parameter of the underlying Einstein-Maxwell Lagrangian, there is no reason to expect it to carry the same value across all BHs. In the language of Ref. Vagnozzi et al. (2023), the regularizing parameter for all three PRBHs considered is a “specific hair” rather than an “universal hair” (see Ref. Vagnozzi et al. (2023) for various examples of BH solutions carrying universal hair), unless one were able to tie $\ell$ to some fundamental parameter of the underlying theory, which however is not the case in the phenomenological approach we are following. In principle one should therefore account for the (non-monochromatic) $\ell$ distribution for PRBHs across the Universe to determine constraints on $f_{\text{pbh}}$. We see no obvious way of doing this, while noting that such a procedure would most likely result in upper limits on $f_{\text{pbh}}$ lying between the Schwarzschild and extremal cases: this observation suffices for our pilot study, and we defer a more complete investigation to future work.
@@
Our final comment concerns the fact that evaporation limits on the PBH abundance are not the only ones at play. Indeed, as recently summarized in Ref. Carr and Kuhnel (2022), there are essentially four classes of limits, each of which is relevant in a different mass range: evaporation, lensing, dynamical, and accretion constraints. Constraints from the accretion of background gas at early times are relevant in a completely different mass range ($10^{33}\lesssim M_{\text{pbh}}/{\text{g}}\lesssim 10^{40}$ – see Fig. 7 of Ref. Carr and Kuhnel (2022) and Fig. 10 of Ref. Carr et al. (2021d)). Although these have been derived assuming Schwarzschild PBHs, moving to the PRBH picture we have considered will not shift the relevant mass range by the $\gtrsim 18$ orders of magnitude required for these constraints to compete with the evaporation ones, unless the physics of gas accreting around RBHs changes drastically with respect to the standard picture, which appears very unlikely. Dynamical constraints, most of which are associated to the destruction of different astronomical objects by the passage of nearby PBHs, are also relevant in a completely different mass range ($10^{34}\lesssim M_{\text{pbh}}/{\text{g}}\lesssim 10^{55}$ – see Fig. 7 of Ref. Carr and Kuhnel (2022) and Fig. 10 of Ref. Carr et al. (2021d)), and considerations completely analogous to those we made for accretion constraints hold.  [^7] Potential exceptions to this mass range for dynamical constraints are those from capture of PBHs by white dwarfs or neutron stars at the centres of globular clusters Capela et al. (2013a, b); Pani and Loeb (2014), or from supernovae explosions resulting from transit of a PBH through a white dwarf Graham et al. (2015). However, these limits are highly disputed because of uncertainties in the dark matter density in globular clusters Ibata et al. (2013); Defillon et al. (2014), or based on the results of hydrodynamical simulations Montero-Camacho et al. (2019). For these reason, we will not consider the previously mentioned limits in our discussion.
@@
Of potentially more relevance to the present work are lensing constraints, which constrain the abundance of PBHs (and more generally massive compact halo objects) with masses $M_{\text{pbh}}\gtrsim 10^{15}\,{\text{g}}$. Indeed, it is lensing constraints which locate the upper edge of the asteroid mass window where PBHs can make up all the DM. Nevertheless, we expect that these constraints should not change when moving from the Schwarzschild PBH framework to the PRBHs considered in this work. Indeed, with all other quantities being fixed (mass of source, relative distances, and so on), lensing constraints only depend on the lens mass $M$, and are unaffected by the metric structure of the lens. Therefore, at fixed mass $M$, we can assume that the lensing limits on Schwarzschild PBHs hold for our PRBHs as well. Note that, as already pointed out in footnote 2, the parameter $M$ appearing in the RBH metrics can be unambiguously identified with the RBH mass, just as with the parameter $M$ in the Schwarzschild metric. We can therefore conclude that for the PRBHs we are considering it should only be the lower edge of the asteroid mass window which is altered with respect to the Schwarzschild case, but not the upper edge. In other words, space-times for which the lower edge moves towards lower masses (as in the Bardeen, Hayward, and CGSV PRBH cases) genuinely correspond to an enlarged asteroid mass window. Therefore, the window where Bardeen, Hayward, and CGSV PRBHs could account for all the DM is much larger compared to that of Schwarzschild PBHs.
@@ V Conclusions
Over the past decade, primordial black holes have regained tremendous interest as viable dark matter candidates, with the so-called “asteroid mass window” ($10^{17}\,{\text{g}}\lesssim M_{\text{pbh}}\lesssim 10^{23}\,{\text{g}}$) where PBHs could potentially account for the entire DM currently still open. Nearly all works on PBHs assume that these are Schwarzschild or Kerr BHs. However, while phenomenologically perfectly valid, such an assumption may stir some unease on the theoretical side, due to the appearance of singularities in these metrics. In our work, we have conducted a pilot study aimed at addressing a question which naturally merges the DM and singularity problems, arguably two among the most important open problems in theoretical physics: “What if PBHs are non-singular”? Our study of primordial regular BHs (PRBHs) has focused on three so-called tr-symmetric metrics (including the well-known Bardeen and Hayward space-times), whereas our companion paper Calzà et al. (2024a) considers non-tr-symmetric metrics, including various metrics inspired from loop quantum gravity.
@@
We show that evaporation constraints on $f_{\text{pbh}}$, the fraction of DM in the form of PRBHs, can be substantially loosened when moving away from the Schwarzschild picture, leading to the asteroid mass window further opening up. For the three PRBHs considered (the Bardeen, Hayward, and CGSV ones) the lower edge of the asteroid mass window is shifted by a decade in mass or more, leading to a larger region of parameter space where PRBHs could account for the entire DM component, which should be the target of the same probes proposed to test the standard window Ray et al. (2021); Kainulainen et al. (2021); Ghosh and Mishra (2024); Branco et al. (2023); Amaral and Schiappacasse (2024); Bertrand et al. (2023); Tran et al. (2024); Dent et al. (2024); Tamta et al. (2024). The nature of the regular BH core (de Sitter or Minkowski) does not appear to play a significant role in this sense. Overall, we have shown that the phenomenology of primordial regular BHs can be particularly rich, making the associated simultaneous solution to the DM and singularity problems one worthy of further study. We note that the constraints we obtained would become even weaker had we moved closer to the extremal limit for the regularizing parameter $\ell$: indeed, in this limit the temperatures of the RBHs we considered approaches zero, implying that there are no constraints from evaporation Pacheco (2018).
@@
We remark that the present work (alongside our companion paper Calzà et al. (2024a)) should be intended as a pilot study, and there are a huge number of interesting follow-up directions. One interesting avenue for further work involves systematically revisiting other sources of constraints which have been studied in the Schwarzschild PBH case, including but not limited to lensing, accretion, and dynamical constraints: while we have argued that these should not alter our considerations on the asteroid mass window, a detailed study which would allow us to extend our constraints over a much larger region of $M_{\text{pbh}}$-$f_{\text{pbh}}$ plane is nevertheless in order. In addition, the metrics we have considered are inherently phenomenological in nature, and it would therefore be worth extending our study to non-singular metrics which enjoy a strong theoretical motivation (our companion paper Calzà et al. (2024a) goes partially in this direction), including potentially metrics which are coupled to the cosmological expansion Farrah et al. (2023); Cadoni et al. (2023, 2024); Faraoni and Rinaldi (2024); Calzà et al. (2024b). Moreover, the formation mechanisms for these PRBHs is likely to be much more complex than the corresponding mechanisms for Schwarzschild PBHs: in fact, the space-times considered are not vacuum solutions of GR, which implies that Birkhoff’s theorem does not (necessarily) hold, and correspondingly the endpoint of gravitational collapse may not be unique. At this stage it is impossible for us to make definitive statements on the matter without abandoning our agnostic viewpoint and assuming a specific underlying theoretical model, but the issue of PRBH formation (possibly within the context of inflationary models leading to an enhanced spectrum of curvature fluctuations over specific scales) is one which requires further study. Last but definitely not least, if PBHs are truly regular, one would hope to ascertain this via signatures complementary to those we have studied: gravitational wave observations, VLBI imaging, motion around BHs, or energy extraction are potentially interesting observables in this sense. For instance, the shadows of Bardeen, Hayward, and CGSV BHs are smaller than their Schwarzschild counterparts by up to $20\%$ Vagnozzi et al. (2023), whereas the motion of stars around these BHs (take for instance the S2 star orbiting around Sgr A) would be altered compared to the motion around a Schwarzschild BH. These probes are being actively studied within the community as a means of distinguishing Schwarzschild BHs from alternatives thereto, and are likely to provide a promising route towards testing the regular nature of PRBHs, in the event that these are detected in the future and demonstrated to make up the DM, or even just a fraction thereof. We plan to address these and other related points in follow-up work.